OpenStudy (anonymous):

What is the radius of the circle whose equation is x^2 + y^2 = 8? 64?

OpenStudy (anonymous):

standard form of equation of circle with center (a,b) and radius r is \[\Large (x-a)^2+(y-b)^2=r^2\] compare your equation with this to get answer!

OpenStudy (anonymous):

hmmm 4?

OpenStudy (anonymous):

What is square root of 8 ??

OpenStudy (anonymous):

nopes! try again look that by comparing \[\Large r^2=8\]

OpenStudy (anonymous):

2\[\sqrt{2}\]

OpenStudy (anonymous):

YEs...

OpenStudy (anonymous):

Well Done..